OpenStudy (anonymous):

Hannah tells you that when variables are in the denominator, the equation four over seven plus eight over x equals one over two becomes solvable. She explains, "There is a value for x that makes the denominator zero, and you can't divide by zero."

OpenStudy (anonymous):

4/5 + 3/x =1/2

OpenStudy (anonymous):

You can solve it by multiplying by the common denominator. Do you what it is?

OpenStudy (anonymous):

umm 10x?

OpenStudy (anonymous):

Right, so if you multiply all the fractions by 10x they will cancel

OpenStudy (anonymous):

ok i got 8x+30/ 10x =5x/10x

OpenStudy (anonymous):